# Q3R: Quadratic Reweighted Rank Regularizer for Effective Low-Rank Training

### v1

@@ 6 Limitations
Arguably, a limitation of this work is also the fact that while the final weights after training are (for appropriate parameters) low-rank, AdamQ3R still handles dense weight matrix variables throughout training, which does not allow a reduction of the parameter budget during training, unlike recent work [47]. More elaborate post-training postprocessing (e.g., inspired by [60]) might lead to further performance improvements.
@@ 7 Conclusion
We introduced Quadratic Reweighted Rank Regularization (Q3R), a principled, optimizer-compatible framework for inducing low-rank structure in deep neural network weights through explicit, continuous regularization. By majorizing a smoothed log-determinant surrogate with a quadratic model embedded in the AdamQ3R optimizer, Q3R trains weight matrices to achieve target ranks with minimal accuracy loss. This enables model compression with negligible performance degradation under reasonable parameter reductions, decreasing deployment costs and increasing throughput. Our experimental results demonstrate that Q3R generalizes across modalities and training regimes, with its design being particularly suitable for low-rank pre-training. Reducing Q3R’s computational overhead, for example via low-rank subspace projections, remains to future work.
Acknowledgement
@@ Appendix A Relationship between Smoothed Log-Determinant and Q3R
In this section, we expand on the relationship between the $\epsilon$-smoothed log-determinant surrogate objective $F_{\epsilon}(\cdot)$ defined in Equation˜1. Part of this material is covered in [33, Section B.2] in a different context.
A.1 Properties of Smoothed Log-Deteriminant
@@
Let now $d:=\min(d_{1},d_{2})$. More precisely, we can define, following [39, 4, 43], a spectral function $F:\operatorname{\mathbb{R}}^{d_{1}\times d_{2}}\to\operatorname{\mathbb{R}}$ as a function for which there exists a function $f:\operatorname{\mathbb{R}}^{d}\to\operatorname{\mathbb{R}}$ for which $F=f\circ\sigma$, where $\sigma:\operatorname{\mathbb{R}}^{d_{1}\times d_{2}}\to\operatorname{\mathbb{R}}^{d},\mathbf{W}\mapsto\sigma(\mathbf{W})=(\sigma_{1}(\mathbf{W}),\ldots,\sigma_{d}(\mathbf{W}))$ is the function mapping matrices in $\operatorname{\mathbb{R}}^{d_{1}\times d_{2}}$ to its singular value vector $\sigma(\mathbf{W})$. A key towards understanding the derivative structure is that we can obtain an explicit formula for the gradient $\nabla F(\mathbf{W})$ of $F_{\epsilon}$ at $\mathbf{W}$ if the function $f$ in the spectral function definition is absolutely (permutation) symmetric [4, Section 7.3] according to Definition˜A.1. It is easy to check that $f_{\epsilon}$ from the definition of the $\epsilon$-smoothed log-determinant $F_{\epsilon}(\cdot)$ satisfies this definition.
@@
For ease of notation, given a vector $\mathbf{v}\in\operatorname{\mathbb{R}}^{d}$, we define $\operatorname{dg}(\mathbf{v})\in\operatorname{\mathbb{R}}^{d_{1}\times d_{2}}$ be the rectangular diagonal matrix such that for $v\in\operatorname{\mathbb{R}}^{d}$ and any $i\in\{1,\ldots,d_{1}\}$, $j\in\{1,\ldots,d_{2}\}$,
@@
Next, we cite a key result about the differentiability of spectral functions which is due to
Proposition A.1 (Differentiability of Spectral Functions [43, Section 7]).
@@
$\nabla F(\mathbf{W})=\mathbf{U}\operatorname{dg}\big(\nabla f(\sigma(\mathbf{W})\big)\mathbf{V}^{\top}$
@@
Since $\lim_{\sigma\nearrow\epsilon}f_{\epsilon}^{\prime}(\sigma)=\epsilon=\lim_{\sigma\searrow\epsilon}f_{\epsilon}^{\prime}(\sigma)$, it follows that $f_{\epsilon}$ is also differentiable at $\sigma=\epsilon$ with $f_{\epsilon}^{\prime}(\epsilon)=\epsilon$, and thus, differentiable on the entirety of its domain. The formula Equation˜9 follows then directly from Proposition˜A.1. ∎
@@
It is well-known that the optimization of convex functions [7] is from many perspectives less challenging than the optimization of non-convex functions. In Section˜4.1, we have claimed that the $\epsilon$-smoothed log-determinant surrogate is not convex. This can indeed by shown directly by invoking [43, Proposition 6.1], which states that a spectral function $F=f\circ\sigma$ is convex if and only if $f$ is convex. Indeed, it is easy to see that $f_{\epsilon}(\cdot)$ is not convex due to its logarithmic dependence on the input for large inputs, which shows that $F_{\epsilon}(\cdot)$ is not a convex function.
As mentioned in Section˜4.1, we see from Equation˜9 that computing $\nabla F_{\epsilon}(\mathbf{W})$ given the matrix $\mathbf{W}\in\operatorname{\mathbb{R}}^{d_{1}\times d_{2}}$ indeed would require a full singular value decomposition that includes at least $d$ leading singular values. Defining $r(\mathbf{W},\epsilon):=|\{i\in\{1,\ldots,d\}:\sigma_{i}(\mathbf{W})>\epsilon\}|$ as in Equation˜3 and $\Sigma=\operatorname{diag}(\sigma_{i}(\mathbf{W}))_{i=1}^{r(\mathbf{W},\epsilon)}\in\operatorname{\mathbb{R}}^{r(\mathbf{W},\epsilon)\times r(\mathbf{W},\epsilon)}$ as in Equation˜4, we obtain for $\mathbf{U}$ and $\mathbf{V}$ defined from the $r(\mathbf{W},\epsilon)$ leading columns of $\mathbf{U}_{\mathbf{W}}=\begin{bmatrix}\mathbf{U}&&\mathbf{U}_{\perp}\end{bmatrix}\in\operatorname{\mathbb{R}}^{d_{1}\times d_{1}}$ and $\mathbf{V}_{\mathbf{W}}=\begin{bmatrix}\mathbf{V}&&\mathbf{V}_{\perp}\end{bmatrix}\in\operatorname{\mathbb{R}}^{d_{2}\times d_{2}}$ that
@@
inserting the formula from Equation˜9, with the notation that $\Sigma_{\perp}=\operatorname{dg}\left(\sigma_{i}(\mathbf{W})\right)_{i=r(\mathbf{W},\epsilon)+1}^{d}$. Fundamentally, this is the key reason why a direct inclusion of the smoothed log-determinant objective into a gradient-based optimization algorithm is computationally inefficient.
@@ A.2 The Quadratic Model Function Underlying Q3R
with $\widetilde{\sigma}_{i}:=\max(\sigma_{i}(\mathbf{W}^{\prime})/\epsilon,1)^{-1}$ for $i\in\{1,\ldots,\max(d_{1},d_{2})\}$, with using the cyclicity of the trace in the third equality. Since $\widetilde{\sigma}_{i},\widetilde{\sigma}_{j}>0$ for all $i,j$, we can establish the lower bound
$\begin{split}&\langle\mathbf{W},\mathcal{R}_{\mathbf{W}^{\prime},\epsilon}(\mathbf{W})\rangle=\sum_{i=1}^{d_{1}}\sum_{j=1}^{d_{2}}\widetilde{\sigma}_{i}\widetilde{\sigma}_{j}\geq\min_{i=1}^{\max(d_{1},d_{2})}\widetilde{\sigma}_{i}^{2}\sum_{i=1}^{d_{1}}\sum_{j=1}^{d_{2}}\mathbf{Z}_{ij}^{2}=\min_{i=1}^{\max(d_{1},d_{2})}\widetilde{\sigma}_{i}^{2}\|\mathbf{Z}\|_{F}^{2}\\ &=\min_{i=1}^{\max(d_{1},d_{2})}\widetilde{\sigma}_{i}^{2}\|\mathbf{U}^{\prime\top}\mathbf{W}\mathbf{V}^{\prime}\|_{F}^{2}=\min_{i=1}^{\max(d_{1},d_{2})}\widetilde{\sigma}_{i}^{2}\|\mathbf{W}\|_{F}^{2}>0.\end{split}$
@@
With a similar argument as made in Equation˜10, we can see that
@@
where $\mathbf{V}\in\operatorname{\mathbb{R}}^{d_{2}\times r(\mathbf{W}^{\prime},\epsilon)}$. Inserting these two equations into Equation˜11, we obtain
@@
Let $\epsilon>0$ For any $\mathbf{W}\in\operatorname{\mathbb{R}}^{d_{1}\times d_{2}}$, the reweighting operator $\mathcal{R}_{\mathbf{W},\epsilon}:\operatorname{\mathbb{R}}^{d_{1}\times d_{2}}\to\operatorname{\mathbb{R}}^{d_{1}\times d_{2}}$ satisfies
@@
This is this a direct implication of Lemma˜A.4 since
@@
using the self-adjointness (see, e.g., [33, Appendix B]) of $\mathcal{R}_{\mathbf{W},\epsilon}$.
The gradient condition Equation˜12 enables us to equate the definition of the quadratic model function Equation˜2 $Q_{\epsilon}(\cdot\mid\mathbf{W}^{\prime})$ (which is, up to a constant that depends on $\epsilon$ and $\mathbf{W}^{\prime}$ the same as the value of Q3R) with the standard quadratic model form of Equation˜5.
@@
Let $\mathbf{W},\mathbf{W}^{\prime}\in\operatorname{\mathbb{R}}^{d_{1}\times d_{2}}$ be arbitrary. To show the equation Equation˜5, we start with its right hand side. By inserting Equation˜12, we obtain
@@
Appendix B Computation of Q3R value
In this section, we provide an implementable algorithm for evaluating the Q3R regularizer $\operatorname{Q3R}_{\mathbf{W}^{\prime},\epsilon}(\mathbf{W})$ as defined in Equation˜6, defined in Algorithm˜3 below.
@@ Appendix B Computation of Q3R value
First, we decompose the reweighting operator image such that
@@
$I_{1}$ $=\bigl\langle\mathbf{W},\,T_{1}^{\epsilon}\bigr\rangle=\epsilon^{2}\,\mathrm{tr}\!\bigl(\mathbf{W}^{\top}\mathbf{U}\,\mathbf{\Sigma}^{-1}\mathbf{U}^{\top}\mathbf{W}\,\mathbf{V}\,\mathbf{\Sigma}^{-1}\mathbf{V}^{\top}\bigr),$
$I_{2}$ $=\bigl\langle\mathbf{W},\,T_{2}^{\epsilon}\bigr\rangle=\epsilon\,\mathrm{tr}\!\bigl(\mathbf{W}^{\top}\mathbf{U}\,\mathbf{\Sigma}^{-1}\mathbf{U}^{\top}\mathbf{W}\,(\mathbf{I}-\mathbf{V}\mathbf{V}^{\top})\bigr),$
$I_{3}$ $=\bigl\langle\mathbf{W},\,T_{3}^{\epsilon}\bigr\rangle=\epsilon\,\mathrm{tr}\!\bigl(\mathbf{W}^{\top}(\mathbf{I}-\mathbf{U}\mathbf{U}^{\top})\,\mathbf{W}\,\mathbf{V}\,\mathbf{\Sigma}^{-1}\mathbf{V}^{\top}\bigr),$
$I_{4}$ $=\bigl\langle\mathbf{W},\,T_{4}^{\epsilon}\bigr\rangle=\mathrm{tr}\!\bigl(\mathbf{W}^{\top}(\mathbf{I}-\mathbf{U}\mathbf{U}^{\top})\,\mathbf{W}\,(\mathbf{I}-\mathbf{V}\mathbf{V}^{\top})\bigr),$
we can write
$\bigl\langle\mathbf{W},\mathcal{R}_{\mathbf{W}^{\prime},\epsilon}(\mathbf{W})\bigr\rangle=I_{1}+I_{2}+I_{3}+I_{4}.$
Apply cyclicity to $I_{1},I_{2},I_{3}$.

### v2

@@ 6 Limitations
Arguably, a limitation of this work is also the fact that while the final weights after training are (for appropriate parameters) low-rank, AdamQ3R still handles dense weight matrix variables throughout training, which does not allow a reduction of the parameter budget during training, unlike recent work [48]. More elaborate post-training post-processing (e.g., inspired by [59]) might lead to further performance improvements.
@@ 7 Conclusion
We introduced Quadratic Reweighted Rank Regularizer (Q3R), a principled, optimizer-compatible framework for inducing low-rank structure in deep neural network weights through explicit, continuous regularization. By majorizing a smoothed log-determinant surrogate with a quadratic model embedded in the AdamQ3R optimizer, Q3R trains weight matrices to achieve target ranks with minimal accuracy loss. This enables model compression with negligible performance degradation under reasonable parameter reductions, decreasing deployment costs and increasing throughput. Our experimental results demonstrate that Q3R generalizes across modalities and training regimes, with its design being particularly suitable for low-rank pre-training. Reducing Q3R’s computational overhead, for example via low-rank subspace projections, remains to future work.
Acknowledgements
@@ Appendix A Relationship between Smoothed Log-Determinant and Q3R
In this section, we expand on the relationship between the $\epsilon$-smoothed log-determinant surrogate objective $F_{\epsilon}(\cdot)$ defined in Equation˜1. Part of this material is covered in [35, Section B.2] in a different context.
A.1 Properties of Smoothed Log-Determinant
@@
Let now $d:=\min(d_{1},d_{2})$. More precisely, we can define, following [39, 3, 38], a spectral function $F:\operatorname{\mathbb{R}}^{d_{1}\times d_{2}}\to\operatorname{\mathbb{R}}$ as a function for which there exists a function $f:\operatorname{\mathbb{R}}^{d}\to\operatorname{\mathbb{R}}$ for which $F=f\circ\sigma$, where $\sigma:\operatorname{\mathbb{R}}^{d_{1}\times d_{2}}\to\operatorname{\mathbb{R}}^{d},\mathbf{W}\mapsto\sigma(\mathbf{W})=(\sigma_{1}(\mathbf{W}),\ldots,\sigma_{d}(\mathbf{W}))$ is the function mapping matrices in $\operatorname{\mathbb{R}}^{d_{1}\times d_{2}}$ to its singular value vector $\sigma(\mathbf{W})$. A key towards understanding the derivative structure is that we can obtain an explicit formula for the gradient $\nabla F(\mathbf{W})$ of $F_{\epsilon}$ at $\mathbf{W}$ if the function $f$ in the spectral function definition is absolutely (permutation) symmetric [3, Section 7.3] according to Definition˜A.1. It is easy to check that $f_{\epsilon}$ from the definition of the $\epsilon$-smoothed log-determinant $F_{\epsilon}(\cdot)$ satisfies this definition.
@@
For ease of notation, given a vector $\mathbf{v}\in\operatorname{\mathbb{R}}^{d}$, we define $\operatorname{dg}(\mathbf{v})\in\operatorname{\mathbb{R}}^{d_{1}\times d_{2}}$ to be the rectangular diagonal matrix such that for $v\in\operatorname{\mathbb{R}}^{d}$ and any $i\in\{1,\ldots,d_{1}\}$, $j\in\{1,\ldots,d_{2}\}$,
@@
Next, we cite a key result about the differentiability of spectral functions, which is due to [38, Section 7].
Proposition A.1 (Differentiability of Spectral Functions [38, Section 7]).
@@
$\nabla F(\mathbf{W})=\mathbf{U}\operatorname{dg}\big(\nabla f(\sigma(\mathbf{W}))\big)\mathbf{V}^{\top}$
@@
Since $\lim_{\sigma\nearrow\epsilon}f_{\epsilon}^{\prime}(\sigma)=\epsilon=\lim_{\sigma\searrow\epsilon}f_{\epsilon}^{\prime}(\sigma)$, it follows that $f_{\epsilon}$ is also differentiable at $\sigma=\epsilon$ with $f_{\epsilon}^{\prime}(\epsilon)=\epsilon$, and thus, differentiable on the entirety of its domain. The formula Equation˜10 follows then directly from Proposition˜A.1. ∎
@@
It is well-known that the optimization of convex functions [5] is from many perspectives less challenging than the optimization of non-convex functions. In Section˜4.1, we have claimed that the $\epsilon$-smoothed log-determinant surrogate is not convex. This can indeed be shown directly by invoking [38, Proposition 6.1], which states that a spectral function $F=f\circ\sigma$ is convex if and only if $f$ is convex. Indeed, it is easy to see that $f_{\epsilon}(\cdot)$ is not convex due to its logarithmic dependence on the input for large inputs, which shows that $F_{\epsilon}(\cdot)$ is not a convex function.
As mentioned in Section˜4.1, we see from Equation˜10 that computing $\nabla F_{\epsilon}(\mathbf{W})$ given the matrix $\mathbf{W}\in\operatorname{\mathbb{R}}^{d_{1}\times d_{2}}$ indeed would require a full singular value decomposition that includes at least $d$ leading singular values. Defining $r(\mathbf{W},\epsilon):=|\{i\in\{1,\ldots,d\}:\sigma_{i}(\mathbf{W})>\epsilon\}|$ as in Equation˜3 and $\Sigma=\operatorname{diag}(\sigma_{i}(\mathbf{W}))_{i=1}^{r(\mathbf{W},\epsilon)}\in\operatorname{\mathbb{R}}^{r(\mathbf{W},\epsilon)\times r(\mathbf{W},\epsilon)}$ as in Equation˜4, we obtain for $\mathbf{U}$ and $\mathbf{V}$ defined from the $r(\mathbf{W},\epsilon)$ leading columns of $\mathbf{U}_{\mathbf{W}}=\begin{bmatrix}\mathbf{U}&&\mathbf{U}_{\perp}\end{bmatrix}\in\operatorname{\mathbb{R}}^{d_{1}\times d_{1}}$ and $\mathbf{V}_{\mathbf{W}}=\begin{bmatrix}\mathbf{V}&&\mathbf{V}_{\perp}\end{bmatrix}\in\operatorname{\mathbb{R}}^{d_{2}\times d_{2}}$ that
@@
inserting the formula from Equation˜10, with the notation that $\Sigma_{\perp}=\operatorname{dg}\left(\sigma_{i}(\mathbf{W})\right)_{i=r(\mathbf{W},\epsilon)+1}^{d}$. Fundamentally, this is the key reason why a direct inclusion of the smoothed log-determinant objective into a gradient-based optimization algorithm is computationally inefficient.
@@ A.2 The Quadratic Model Function Underlying Q3R
with $\widetilde{\sigma}_{i}:=\max(\sigma_{i}(\mathbf{W}^{\prime})/\epsilon,1)^{-1}$ for $i\in\{1,\ldots,\max(d_{1},d_{2})\}$, using the cyclicity of the trace in the third equality. Since $\widetilde{\sigma}_{i},\widetilde{\sigma}_{j}>0$ for all $i,j$ (with convention that $\sigma_{i}(\mathbf{W}^{\prime})=0$ for $i>\min(d_{1},d_{2})$), we can establish the lower bound
$\begin{split}&\langle\mathbf{W},\mathcal{R}_{\mathbf{W}^{\prime},\epsilon}(\mathbf{W})\rangle=\sum_{i=1}^{d_{1}}\sum_{j=1}^{d_{2}}\widetilde{\sigma}_{i}\,\widetilde{\sigma}_{j}\,\mathbf{Z}_{ij}^{2}\geq\min_{i=1}^{\max(d_{1},d_{2})}\widetilde{\sigma}_{i}^{2}\sum_{i=1}^{d_{1}}\sum_{j=1}^{d_{2}}\mathbf{Z}_{ij}^{2}=\min_{i=1}^{\max(d_{1},d_{2})}\widetilde{\sigma}_{i}^{2}\|\mathbf{Z}\|_{F}^{2}\\ &=\min_{i=1}^{\max(d_{1},d_{2})}\widetilde{\sigma}_{i}^{2}\|\mathbf{U}^{\prime\top}\mathbf{W}\mathbf{V}^{\prime}\|_{F}^{2}=\min_{i=1}^{\max(d_{1},d_{2})}\widetilde{\sigma}_{i}^{2}\|\mathbf{W}\|_{F}^{2}>0.\end{split}$
@@
With a similar argument as made in Equation˜11, we can see that
@@
where $\mathbf{V}\in\operatorname{\mathbb{R}}^{d_{2}\times r(\mathbf{W}^{\prime},\epsilon)}$. Inserting these two equations into Equation˜12, we obtain
@@
Let $\epsilon>0$. For any $\mathbf{W}\in\operatorname{\mathbb{R}}^{d_{1}\times d_{2}}$, the reweighting operator $\mathcal{R}_{\mathbf{W},\epsilon}:\operatorname{\mathbb{R}}^{d_{1}\times d_{2}}\to\operatorname{\mathbb{R}}^{d_{1}\times d_{2}}$ satisfies
@@
This is a direct implication of Lemma˜A.4 since
@@
using the self-adjointness (see, e.g., [35, Appendix B]) of $\mathcal{R}_{\mathbf{W},\epsilon}$.
The gradient condition Equation˜13 enables us to equate the definition of the quadratic model function Equation˜2 $Q_{\epsilon}(\cdot\mid\mathbf{W}^{\prime})$ (which is, up to a constant that depends on $\epsilon$ and $\mathbf{W}^{\prime}$ the same as the value of Q3R) with the standard quadratic model form of Equation˜5.
@@
Let $\mathbf{W},\mathbf{W}^{\prime}\in\operatorname{\mathbb{R}}^{d_{1}\times d_{2}}$ be arbitrary. To show the equation Equation˜5, we start with its right hand side. By inserting Equation˜13, we obtain
@@
Appendix B Computation of Regularizer Value of Q3R
In this section, we provide with Algorithm˜3 below an implementable algorithm for evaluating the Q3R regularizer $\operatorname{Q3R}_{\mathbf{W}^{\prime},\epsilon}(\mathbf{W})$ of a weight matrix $\mathbf{W}$ given, as defined in Equation˜6, given the reweighting operator $\mathcal{R}_{\mathbf{W}^{\prime},\epsilon}(\cdot)$ of Definition˜4.1 defined for the matrix $\mathbf{W}^{\prime}$ with leading left and right singular vector matrices $\mathbf{U}\in\operatorname{\mathbb{R}}^{d_{1}\times r}$ and $\mathbf{V}\in\operatorname{\mathbb{R}}^{d_{2}\times r}$ and leading $r$ singular vectors $\sigma\in\operatorname{\mathbb{R}}^{r}$, where $r=r(\mathbf{W}^{\prime},\epsilon)$ as in (3).
@@ Appendix B Computation of Regularizer Value of Q3R
First, we decompose, following its definition (4), the reweighting operator image such that
@@
$I_{1}$ $=\bigl\langle\mathbf{W},\,T_{1}^{\epsilon}\bigr\rangle=\epsilon^{2}\,\operatorname{tr}\bigl(\mathbf{W}^{\top}\mathbf{U}\,\mathbf{\Sigma}^{-1}\mathbf{U}^{\top}\mathbf{W}\,\mathbf{V}\,\mathbf{\Sigma}^{-1}\mathbf{V}^{\top}\bigr),$
$I_{2}$ $=\bigl\langle\mathbf{W},\,T_{2}^{\epsilon}\bigr\rangle=\epsilon\,\operatorname{tr}\bigl(\mathbf{W}^{\top}\mathbf{U}\,\mathbf{\Sigma}^{-1}\mathbf{U}^{\top}\mathbf{W}\,(\mathbf{I}-\mathbf{V}\mathbf{V}^{\top})\bigr),$
$I_{3}$ $=\bigl\langle\mathbf{W},\,T_{3}^{\epsilon}\bigr\rangle=\epsilon\,\operatorname{tr}\bigl(\mathbf{W}^{\top}(\mathbf{I}-\mathbf{U}\mathbf{U}^{\top})\,\mathbf{W}\,\mathbf{V}\,\mathbf{\Sigma}^{-1}\mathbf{V}^{\top}\bigr),$
$I_{4}$ $=\bigl\langle\mathbf{W},\,T_{4}^{\epsilon}\bigr\rangle=\operatorname{tr}\bigl(\mathbf{W}^{\top}(\mathbf{I}-\mathbf{U}\mathbf{U}^{\top})\,\mathbf{W}\,(\mathbf{I}-\mathbf{V}\mathbf{V}^{\top})\bigr),$
and recalling the definition (6) of $\operatorname{Q3R}_{\mathbf{W}^{\prime},\epsilon}(\mathbf{W})$, we can write
$\operatorname{Q3R}_{\mathbf{W}^{\prime},\epsilon}(\mathbf{W})=\frac{1}{2}\bigl\langle\mathbf{W},\mathcal{R}_{\mathbf{W}^{\prime},\epsilon}(\mathbf{W})\bigr\rangle=\frac{1}{2}\left(I_{1}+I_{2}+I_{3}+I_{4}\right).$
Below, we outline several steps that lead to an efficient implementation of the computation of the value of $\operatorname{Q3R}_{\mathbf{W}^{\prime},\epsilon}(\mathbf{W})$. Apply cyclicity to $I_{1},I_{2},I_{3}$.